OpenStudy (anonymous):

If you stand 20 m away from a building , your line of sight to the top of the building is at a 53 degree angle above the horizontal. Your eyes are 1.4 m above the ground. How tall is the building?

OpenStudy (anonymous):

Hi, draw the situation.  When you do so, you will see, you can use tangens function, to create equation as follows: \[\tan 53 = \frac{ x+1,4 }{ 20 }\] Now just compute the tangens and separate x: \[x +1,4 = 20 * 1,32\] and then \[x = 25,14 m\]